OpenStudy (anonymous):

Write the standard equation for the circle. center (–6, 9), r = 3

OpenStudy (anonymous):

y=(x-h)^2+(y-k)^2=r^2 r=radius (h,k)=vertex can you plug those in?

OpenStudy (anonymous):

not vertex, center lolol

OpenStudy (anonymous):

lol yea hold on

OpenStudy (anonymous):

y=(-6-h)^2+(9-k)^2=3^2

hartnn (hartnn):

uh oh.... loose the 'y=' part and h=-6,k=9,r=3 in \(\large (x-h)^2+(y-k)^2=r^2\)